OpenStudy (anonymous):

what is the reasonable estimate of the average kinetic energy of an athlete during a 100m race that takes 10s ? (please show the working aswell. hurry its my exam tomorrow !!! )

OpenStudy (jennychan12):

the speed would be 100/10 = 10 m/x so if that were the velocity, it would be KE = m(10^2). let's assume this athlete's m is 50kg. so it'd be KE = 50(100) so around 5000 J